OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. x2 + 3x - 18 = 0 A. {-6, 3} B. {-2, 9} C. {-3, 6} D. {-9, 2}

OpenStudy (saifoo.khan):

A for apple

hero (hero):

x^2 + 3x - 18 = 0 x^2+6x-3x - 18 = 0 x(x+6)-3(x+6) = 0 (x+6)(x-3) =0 x = -6,3 Rule of thumb: Don't use the quadratic formula if you don't have to.

OpenStudy (anonymous):

\[x=\frac{-3 \pm \sqrt{3^2-4*1*(-18)}}{2(1)}\]\[x=\frac{-3 \pm \sqrt{9+72}}{2}\]\[x=\frac{-3 \pm \sqrt{81}}{2}\]\[x=\frac{-3-9}{2}=-6\]or\[x=\frac{-3+9}{2}=3\]And as saifoo and hero already said the answer is A

OpenStudy (anonymous):

Rule of pinky: follow directions

hero (hero):

Sorry, I'm just doing what I was taught. My teacher said use the quadratic formula as last resort.

OpenStudy (anonymous):

personally, i hate the quadratic formula!

hero (hero):

I hate it worse than you do. I don't even use it anymore.

hero (hero):

I'll never use it again ever in my life.

OpenStudy (anonymous):

ok, you are bigger anti-quadratic-formula-ite

OpenStudy (jamesj):

@bates12: you've been handed the answer on a silver platter. But let me ask you: do you know what the quadratic equation is?